# Emergence of a High-Dimensional Abstraction Phase in Language Transformers

### v1

@@ 4.2 Language processing during the high-dimensionality phase
We just saw, via geometric evidence ($\Delta$), that the high-ID phase marks a change in processing function. Now, we examine exactly what that function is. To do so, we look at layer-wise classification accuracies for the probing tasks described in Appendix H. We find that, in general, ID-peak representations fail at surface-form tasks but excel at semantic and syntactic tasks, indicating a functional transition from superficial to abstract linguistic processing.
@@
In Figure 5(a), we consider the accuracy for two tasks, Sentence Length and Word Content, which test whether a layer retains information about superficial properties of the input. We observe that the ability to correctly reconstruct the length, measured in tokens, of the input sentence gets lost as we climb the network layers. Intriguingly, for OPT (Figure 5(a), center), which is the model with the latest ID peak, initial accuracy is relatively stable, and only starts to significantly decrease at the onset of the peak.
Concerning the Word Content task, which tests the ability to detect the presence of specific words in the input, we generally observe a great decrease, particularly clear during ID expansion. Interestingly, accuracy tends to go up again after the ID peak, probably because, as the model prepares to predict the output, more concrete lexical information is again encoded in its representations. Together, the surface-form tasks confirm the evidence from $\Delta(l_{i}\to l_{first})$ (Figure 2 above) that the ID peak processes a more abstract type of information that, as we are about to see, may relate to the syntactic and semantic contents of the sequence.
@@
Figure 5(b) reports accuracy for Llama, Pythia, and OPT (results for the remaining models in Appendix H) on the syntactic and semantic probe tasks. Despite variation across tasks and models, we observe that, in general, asymptotic accuracy is reached during the ID peak phase. Again, this implies that, for OPT, the asymptote is reached later. For OLMo (Appendix H), we observe in some cases a late accuracy peak, related to the second ID expansion phase that this model undergoes. In general, however, once top accuracy is reached, performance stays quite constant across layers, suggesting that the expressivity of high dimensionality permits rich linguistic representaiton of the inputs, but, once this information is extracted, it is propagated across subsequent layers. This is intuitive, as high-level linguistic information is useful to the network for its ultimate task of next-token prediction.
@@
Given the apparent linguistic importance of the high-ID phase, a natural question concerns the extent to which the nature of the ID peak explains the LM’s performance on its original task of next-token prediction. The question was already partially answered by Figures 1 and D.1, which show an absence of peaks, respectively, when trained LMs process shuffled text and when untrained LMs process sane (unshuffled) text; in both cases, next-token prediction is impossible.
We further computed Spearman correlations between average prediction surprisal for each (model, corpus) combination and the corresponding maximum ID values (Figure 6, left), as well as ID-peak onsets (Figure 6, right). As the plots show, even when limiting the analysis to sane text, where the differences in surprisal will be smaller, there is a marginal tendency for maximum ID value to inversely correlate with surprisal: the higher the peak, the better the LM is at predicting the next token. There is, moreover, a significant positive correlation between surprisal and the onset of the ID peak: the earlier the peak, the better the model is at predicting the next token.
We just saw that the ID peak marks the phase where the model first completes a full syntactic and semantic analysis of the input. The earlier this analysis takes place, the more layers the model will have to further refine its prediction by relying on it. The correlation between ID peak onset and surprisal thus indirectly confirms the importance of the high-dimensional processing phase for good model performance.
Figure 6: Surprisal plotted against the maximum ID across layers (left) and the relative ID peak onset over layers (right), where each datapoint is a (model, corpus) combination ($N=50$k sequences per corpus). A linear fit (left) and log-linear fit (right) are shown. (Left) Surprisal negatively correlates to maximum ID with Spearman $\rho=-0.46$, $p=0.09$, meaning that higher ID indicates better LM performance. (Right) Surprisal positively correlates to ID peak onset, $\rho=0.65$, $p=0.01$, meaning that an earlier ID peak indicates better LM performance.
@@
Finally, given that the ID peak marks an abstract semantic representation of the input, representations at this peak should also viably transfer to downstream semantic tasks. We confirm this hypothesis for sentiment and toxicity classification.
Validation accuracy curves for both tasks are shown for Llama, Pythia and OPT in Figure 7, with more results in Appendix I. Similar to the semantic and syntactic probing results, downstream classification performance consistently converges at the ID peak across models and remains high thereafter. Note that, while the ID peak is computed from the generic input corpora Bookcorpus, the Pile, and Wikitext, it still predicts transferrability to different downstream datasets that are reasonably in-distribution.
@@ 5 Conclusion
It is evident that a LM needs to extract information from its input to predict the next token. The non-trivial fact we show here is that this process does not gradually refine representations, but rather undergoes a phase transition characterized by ID expansion, cross-model information sharing, and a switch to abstract information processing. Mirroring Jastrzebski et al. (2018)’s proposition for visual networks, our findings are compatible with a view of LMs in which the high-dimensional early-to-mid layers functionally specialize to analyze inputs in a relatively fixed manner, whereas later layers may more flexibly refine the output prediction, using information extracted during the high-ID phase. Interestingly, two recent papers (Gromov et al., 2024; Men et al., 2024) found that (1) late LM layers better approximate each other than earlier layers do, and (2) pruning late layers (excluding the last) affects performance less than pruning earlier layers. This fully aligns with our results and suggests a need to test the effect of pruning inside and outside the ID peak. Other studies have highlighted the importance of central layers in performing various core functions. For example, Hendel et al. (2023) find that compositional “task vectors” are formed in layers superficially consistent with the peaks we detect. Again, future work should more thoroughly study the relation between the ID profile and specific circuits detected in mechanistic interpretability work.
@@
While our work closely relates to that of Valeriani et al. (2023), who studied the evolution of ID in vision and protein transformers, an interesting contrast is that they found crucial semantic information to coalesce during a dimensionality reduction phase, whereas we associated similar marks to a dimensionality expansion phase. In line with our results, recent work in theoretical neuroscience shows that high ID, thanks to its expressivity, underlies successful few-shot learning (Sorscher et al., 2022) and generalizable latent representations for DNNs (Elmoznino and Bonner, 2024; Wakhloo et al., 2024). Conversely, primarily in artificial and biological vision, low ID is linked to generalization thanks to representations’ robustness to noise and greater linear separability in embedding space (Amsaleg et al., 2017; Chung et al., 2018; Cohen et al., 2020). Clearly, whether dimensionality is a curse or blessing to performance depends on the context of the learning problem. Reconciling, then, why and when high performance arises from reduced or expanded dimensionality remains an important direction for future work.
6 Limitations
  • While we establish an empirical connection between the ID peak and linguistic processing, we do not offer a mechanistic understanding of how the latter results in ID expansion, including which differences between models lead to different patterns, such as an earlier or later ID peak onset.
  • The OLMo model displays a second ID peak and various outlying patterns that we did not attempt to interpret.
  • We explored 5 language models on data extracted from three different corpora. However, due to compute constraints, the models we picked are all in the same size range (between 6 and 8 billion parameters). In the future, we would like to explore the extent to which the patterns we detected also appear at other scales.
@@ C.1 Scale analysis
For each model and corpus, we perform a scale analysis to select the neighbor order. An example of such an analysis is shown in Figure C.1, performed on the Pile for Pythia. We plot the ID estimate for increasing scales (x-axis in Figure C.1) for each layer. The true ID is likely to lie at a scale in which the ID estimate plateaus [Denti et al., 2022], which is marked by the highlighted region. For simplicity, per model-corpus combination, we choose one scale for all layers: in this particular example, we choose order $k=2^{5}$. In general, the scale tends to be around $k=32$ for all (model, corpus) combinations, allowing us to reliably compare between them (see Tables 1(a) and 1(b) for all $k$). Once the scale is chosen for each model-corpus combination, we plot the scale-adjusted ID estimates (see, e.g., Figure C.2).
Figure C.1: GRIDE ID estimation neighborhood scale analysis example for Pythia on the Pile on a single random seed, where each line is a layer’s ID estimate at different scales. All layers shown reach a plateau in the highlighted range.
@@
(a) GRIDE order $k$ reported for each model, corpus, mode (in sane and shuffled) combination. For simplicity, we chose one $k$ for all layers.
@@ C.2 Additional results
Figure C.2 displays the evolution of estimated ID, scale-adjusted, per layer for all corpora and models. While the magnitude of ID differs across corpora, with ID on Bookcorpus lower than those of the Pile and Wikitext for all models, all corpora exhibit the characteristic high-ID peak, and at nearly the same onset. The dampened Bookcorpus peak IDs, which are still significantly above the corresponding shuffled ID peaks, might be explained by the fact that this corpus is entirely made of novels, and it is thus the less in-domain of the corpora we used.
@@
Appendix D Intrinsic dimension over training
Figure D.1 shows the evolution of the average-ID profile over the course of training for Pythia, whose checkpoints are publicly available. Of-note, the ID profile seems to converge at between 16000 and 64000 iterations, or less than halfway until training is finished. We observe that towards the beginning of training (blue line), the ID profile is quite flat, and that, over the course of training, the characteristic central ID peak emerges. Beyond shape, we see that, while the initial-layer ID remains the same across checkpoints, the IDs of other layers tend to increase in magnitude. Both observations suggest that expansion of ID underlies linguistic processing of the input.
Figure D.1: Evolution of the average-ID profile over training for Pythia. Each curve corresponds to one checkpoint, at steps 512 until 143000 (the final model), and is shown with $\pm$ 2 SD over corpora and partitions.
@@ Appendix G Cross-model $\Delta$
Appendix H Probing tasks
H.1 Tasks
We use the following classification tasks from Conneau et al. [2018]:

### v2

@@ 4.2 Language processing during the high-dimensionality phase
We just saw, via geometric evidence ($\Delta$), that the high-ID phase marks a change in processing function. Now, we attempt to interpret what that function is. To do so, we look at layer-wise classification accuracies for the probing tasks described in Appendix I. We find that, in general, ID-peak representations fail at surface-form tasks but excel at semantic and syntactic tasks, indicating a functional transition from superficial to abstract linguistic processing.
@@
In Figure 5(a), we consider the accuracy for two tasks, Sentence Length and Word Content, which test whether a layer retains information about superficial properties of the input. We observe that the ability to correctly reconstruct the length, measured in tokens, of the input sentence gets lost as we climb the network layers, and performance starts a sharp decrease at the onset of the ID peak or shortly before it. Intriguingly, for OPT (Figure 5(a), center), which is the model with the latest ID peak, initial accuracy is relatively stable, and only starts to significantly decrease at the onset of the peak, further suggesting that it is only during the high-ID phase, even when the latter occurs relatively late, that superficial information (such as the number of tokens in the input sentence) is discarded by the models, that, as we are about to see, start instead at that point to process more abstract syntactic and semantic information.
Concerning the Word Content task, which tests the ability to detect the presence of specific words in the input, we generally observe a great decrease after the first few layers, particularly clear during ID expansion. Interestingly, accuracy tends to go up again after the ID peak, probably because, as the model prepares to predict the output, more concrete lexical information is again encoded in its representations. Together, the surface-form tasks confirm the evidence from $\Delta(l_{i}\to l_{first})$ (Figure 2 above) that the ID peak processes a more abstract type of information that, as we are about to see, may relate to the syntactic and semantic contents of the sequence.
@@
Figure 5(b) reports accuracy for Llama, Pythia, and OPT (results for the remaining models in Appendix I) on the syntactic and semantic probe tasks. As described in detail in Appendix I, these tasks require a model to detect whether a sentence is syntactically well-formed (Bigram Shift), whether it is semantically coherent (Odd Man Out) and whether it describes a pair of events or states in a plausible order (Coordination Inversion). Despite variation across tasks and models, we observe that, in general, asymptotic accuracy is reached during the ID peak phase. Again, this implies that, for OPT, the asymptote is reached later. For OLMo (Appendix I), we observe in some cases a late accuracy peak, related to the second ID expansion phase that this model undergoes. In general, however, once top accuracy is reached, performance stays quite constant across layers, suggesting that the expressivity of high dimensionality permits rich linguistic representaiton of the inputs, but, once this information is extracted, it is propagated across subsequent layers. This is intuitive, as high-level linguistic information is useful to the network for its ultimate task of next-token prediction. Appendix D confirms the same patterns for a smaller and a larger Pythia model.
@@
Given the apparent linguistic importance of the high-ID phase, a natural question concerns the extent to which the nature of the ID peak relates to the LM’s performance on its original task of next-token prediction. The question was already partially answered by Figure 1, which shows an absence of peaks, respectively, when trained LMs process shuffled text and when untrained LMs process normal text; in both cases, next-token prediction is impossible.
We further computed Spearman correlations between average prediction surprisal for each (model, corpus) combination and the corresponding maximum ID values (Figure 6, left), as well as ID-peak onsets (Figure 6, right). [^4] Due to its range, it’s easier to visualize surprisal, instead of the more commonly used perplexity measure, in the plots. As perplexity is exponentiated surprisal, and we are using Spearman correlation (which is robust to monotonic transformations such as exponentiation), our results are not affected by this choice.
As the plots show, even when limiting the analysis to sane text processed by fully trained models, where the differences in surprisal will be smaller, there is a marginal tendency for maximum ID value to inversely correlate with surprisal: the higher the peak, the better the LM is at predicting the next token. There is, moreover, a significant positive correlation between surprisal and the onset of the ID peak: the earlier the peak, the better the model is at predicting the next token.
We just saw that the ID peak might mark the phase where the model first completes a full syntactic and semantic analysis of the input. The earlier this analysis takes place, the more layers the model will have to further refine its prediction by relying on it. The correlation between ID peak onset and surprisal thus indirectly confirms the importance of the high-dimensional processing phase for good model performance.
Figure 6: Surprisal plotted against the maximum ID across layers (left) and the relative ID peak onset over layers (right), where each datapoint is a (model, corpus) combination ($N=50$k sequences per corpus). A linear fit (left) and log-linear fit (right) are shown. (Left): Surprisal negatively correlates to maximum ID with Spearman $\rho=-0.46$, $p=0.09$, meaning that higher ID indicates better LM performance. (Right): Surprisal positively correlates to ID peak onset, $\rho=0.65$, $p=0.01$, meaning that an earlier ID peak indicates better LM performance.
@@
Finally, given that the ID peak seems to mark the onset of more abstract linguistic processing of the input, representations at this peak should also viably transfer to downstream linguistic tasks. We confirm this hypothesis for sentiment and toxicity classification.
Validation accuracy curves for both tasks are shown for Llama, Pythia and OPT in Figure 7, with more results in Appendix J. Similar to the semantic and syntactic probing results, downstream classification performance consistently converges at the ID peak across models and remains high thereafter. Note that, while the ID peak is computed from three generic input corpora (Bookcorpus, Pile and Wikitext), it still predicts transferability to different downstream datasets that are reasonably in-distribution.
@@ 5 Conclusion
It is evident that a LM needs to extract information from its input to predict the next token. The non-trivial fact we show here is that this process does not gradually refine representations, but rather undergoes a phase transition characterized by ID expansion, cross-model information sharing, and a switch to abstract information processing. Our main take-home message is that different LMs consistently develop a central-layer phase where the intrinsic dimension is expanded, which is the locus of deeper linguistic processing.
Our paper focused on detecting broad qualitative patterns. As it is the first study of its kind, we wanted to establish some basic high-level results. Without the expectations set by our high-level investigation, we cannot make precise predictions to be tested in more specialized mechanistic-interpretability studies. We note moreover that, while high-level, our conclusions are supported by very consistent converging evidence from a variety of models, corpora and experiments (even OLMo, that is to some extent the “outlier” model, displays the ID peak and the relevant associated properties). Future work should establish a clearer causal connection between intrinsic dimension and language processing, via layer ablations studies and by constraining the dimensionality of different layers during language processing.
@@
Mirroring Jastrzebski et al. (2018)’s proposition for visual networks, our findings are compatible with a view of LMs in which the high-dimensional early-to-mid layers functionally specialize to analyze inputs in a relatively fixed manner, whereas later layers may more flexibly refine the output prediction, using information extracted during the high-ID phase. Interestingly, two recent papers (Gromov et al., 2024; Men et al., 2024) found that (1) late LM layers better approximate each other than earlier layers do, and (2) pruning late layers (excluding the last) affects performance less than pruning earlier layers. This fully aligns with our results, and suggests a need to test the effect of pruning inside and outside the ID peak. Other studies have highlighted the importance of central layers in performing various core functions. For example, Hendel et al. (2023) find that compositional “task vectors” are formed in layers superficially consistent with the peaks we detect. Again, future work should more thoroughly study the relation between the ID profile and specific circuits detected in mechanistic interpretability work.
@@
While our work closely relates to that of Valeriani et al. (2023), who studied the evolution of ID in vision and protein transformers, an interesting contrast is that they found crucial semantic information to coalesce during a dimensionality reduction phase, whereas we associated similar marks to a dimensionality expansion phase. This difference in observations may be partly attributed to our differing methodologies, since Valeriani and colleagues analyzed the ID of average sequence token and probed semantic information using a classifier based on nearest neighbors. In line with our results, recent work in theoretical neuroscience shows that high ID, thanks to its expressivity, underlies successful few-shot learning (Sorscher et al., 2022) and generalizable latent representations for DNNs (Elmoznino & Bonner, 2024; Wakhloo et al., 2024). Conversely, primarily in artificial and biological vision, low ID is linked to generalization thanks to representations’ robustness to noise and greater linear separability in embedding space (Amsaleg et al., 2017; Chung et al., 2018; Cohen et al., 2020). Clearly, whether dimensionality is a curse or blessing to performance depends on the context of the learning problem. Reconciling, then, why and when high performance arises from reduced or expanded dimensionality remains an important direction for future work.
From a more applied perspective, we see various ways in which our discovery of a consistent high-ID phase could be useful. ID profiles emerge as significant blueprints of model behaviour that could be used as proxies of model quality. ID information can be used for model pruning, or to choose which layers to fine-tune, or for model stitching and other model-interfacing operations. These are all exciting directions for future work.
6 Reproducibility statement
We uploaded corpora, code and a readme document to reproduce our experiments as supplementary materials, and we will make them publicly available upon acceptance. We report our compute usage in Appendix A, and describe all publicly available resources we used in Appendix B. Appendices Appendix C, Appendix I and Appendix J specify relevant hyperparameter choices for ID computation, probing task and transfer task implementation, respectively.
@@ C.1 Scale analysis
In this section we explicitly show how to perform a scale analysis in order to select a proper and meaningful scale when computing the ID. In fact, as explained in the main text, the ID estimation might be affected by undesirable effects that are always present when dealing with real-world datasets and that can hide the true dimensionality of the manifold underlying the data. Such effects include the presence of noise, which typically affects small scales, and density variations and manifold curvature, which, instead, tend to be observed at a larger scale (Denti et al., 2022). It is thus recommended to choose an intermediate scale, where the ID estimate might be stable across different neighbourhood sizes, and it is less likely to be affected by the aforementioned spurious factors. For this reason, it is necessary to see how the ID changes when varying the neighbourhood size taken into account when performing the ID computation. We rely on the GRIDE estimator, which allows to explicitly select the number of neighbours considered. In particular, the rank of the first nearest neighbour used to compute the distance ratio is a hyperparameter, and we refer to the chosen value for this hyperparameter as the scale in what follows.
We performed scale analysis for each model and corpus. An example is provided in Figure C.1, which shows the results for Pythia on Pile. We plot the ID estimate for increasing scales (i.e., number of neighbours, x-axis in Figure C.1) for some representative layers. The rank of the neighbour considered is explored by means of powers of 2 in order to look at regions where the variation of the ID would be noticeable. The true ID is likely to lie at a scale in which the ID estimate approximately plateaus (Denti et al., 2022), which is marked by the highlighted region. For simplicity, per model-corpus combination, we choose one scale for all layers: in this particular example, we choose $k=2^{5}$. In general, the optimal neighbourhood size tends to be around $k=32$ for all (model, corpus) combinations, allowing us to reliably compare them (see Table C.1 for all $k$). Once the scale is chosen for each (model, corpus) combination, we plot the scale-adjusted ID estimates (see, e.g., Figure C.3).
Figure C.1: GRIDE ID estimation neighbourhood scale analysis example for Pythia on the Pile on a single random seed, where each line is a layer’s ID estimate at different scales. All layers shown reach a plateau in the highlighted range.
As a further robustness test, we show in Figure C.2 that results are fairly stable to choosing scales within the shaded range. For each model, we include a panel with the original chosen scale (in blue), e.g., $k=2^{5}=$32th nearest neighbor, and then look at the shape of the ID if we move the scale one (logarithmic) step down $k=2^{4}=$16 (orange) or up $k=2^{6}=$64 (green). Fortunately, we can see that the shapes of the resulting ID curves are stable across all models, and the ID values themselves change very little.
Figure C.2: Robustness of the scale analysis within the plateau region across models. The ID estimates are only weakly affected by doubling (green) or halving (orange) the scale with respect to the results reported in the main text (blue).
@@
(a) GRIDE scale $k$ reported for each model, corpus, mode (in sane and shuffled) combination. For simplicity, we chose one $k$ for all layers.
@@
corpus Pythia size GRIDE $k$
wikitext 2.8b 32
12b 16
(c) GRIDE $k$ for additional Pythia sizes, on Wikitext dataset.
Table C.1: Gride IDs for each model, dataset, and mode
@@ C.2 Additional results
Figure C.3 displays the evolution of estimated ID, scale-adjusted, per layer for all corpora and models. While the magnitude of ID differs across corpora, with ID on Bookcorpus lower than those of the Pile and Wikitext for all models, all corpora exhibit the characteristic high-ID peak, and at nearly the same onset. The dampened Bookcorpus peak IDs, which are still significantly above the corresponding shuffled ID peaks, might be explained by the fact that this corpus is entirely made of novels, and it is thus the less in-domain of the corpora we explored.
@@
Appendix D Influence of model size on ID and probing tasks
We reproduce key experiments from the paper with a smaller (2.8b) and a larger (12b) version of the Pythia’s architecture, suggesting that our findings are not limited to the scale of the models we analyzed in the paper.
Figure D.1: ID profiles (a) and probing-task experiments (b, c) reproduced with a smaller and a larger model from the Pythia family.
ID profiles
We have plotted the ID profiles of smaller (2.8b) and larger (12b) Pythia models in Figure D.1-(a). Analogously to the evolution of ID during training (cf. Figure 1 right), we observe a similar profile across model sizes, with the ID magnitude of the peak growing with model size. Note that the 2.8b and 6.9b model layers share the same ambient dimensionality (4096), whereas the 12b model has an ambient dimensionality of 5120. However, the change in peak magnitude is much more dramatic between 2.8b and 6.9b than between 6.9b and 12b, confirming the observation by Cheng et al. (2023) that ID is only weakly linked to ambient dimensionality.
Probing tasks
We also reproduce the probing-task experiments with the same smaller (2.8b) and larger (12.b) models from the Pythia family, showing that the semantic/syntactic probing-task asymptote is reached within the span of the ID peak also at these newly tested sizes (see Figure D.1-(b,c)).
@@ Appendix E Information imbalance with respect to first/last layer
Figure E.2 shows, in the spirit of the Information Plane (Shwartz-Ziv & Tishby, 2017), the $\Delta(l_{i}\to l_{last})$ plotted against $\Delta(l_{first}\to l_{i})$ for all models across the layers. The trajectory reveals the dynamic evolution of input and output informativity along the residual stream.
@@
Figure E.2: Information Plane: $\Delta(l_{i}\to l_{last})$ vs. $\Delta(l_{first}\to l_{i})$. For each LM, the information imbalance $\Delta$ from the $i$th to the last layer is plotted against $\Delta$ from the first to the $i$th layer, where lower $\Delta$ means more informative. Each point on the curve corresponds to a single layer $i$; for all models, the first layer (gold star) begins on the y-axis, $\Delta(l_{first}\to l_{i})=0$, and the trajectory through the layers ends on the x-axis, $\Delta(l_{i}\to l_{last})=0$. Consistently across models, ID peak layers (red) mark a changepoint in processing patterns. Pre-peak layers show a rapid increase along the x-axis but slow decrease along the y-axis, indicating a rapid departure from the inputs while not learning much about the outputs. This reflects the phase of semantic abstraction. Post-peak layers’ information dynamics vary by model, though all are marked by a final rapid decrease in $\Delta(l_{i}\to l_{last})$ and a decrease in $\Delta(l_{first}\to l_{i})$. This suggests a return to surface-level information processing in order to predict the next token.
@@ Appendix G Cross-model $\Delta$
Appendix H Cross-model layer comparison using CKA
We reproduce the cross-model layer comparison experiments of Figure 4) using linear CKA (Kornblith et al., 2019) as an alternative measure of similarity. Figure H.1 broadly confirms the results found with $\Delta{}$.
(a) Pythia vs. Llama
(b) OPT vs. Llama
(c) Pythia vs. OPT
Figure H.1: Cross-model layer similarity measured using linear CKA. Like in Figure 4, ID-peak intersections tend to coincide with high-similarity areas.
Appendix I Probing tasks
I.1 Tasks
We use the following classification tasks from Conneau et al. (2018):